# Photoacoustic Iterative Optimization Algorithm with Shape Prior Regularization

### v1

@@ 3.2 Vivo Animal Experiment Results
Then in vivo animal study data, including human arm was acquired by Shuang Li’s work[22] using synthetic matrix array. The system employed a nonfocusing linear array (customized by Imasonics, France) to receive PA signals. The linear array has 256 elements with a pitch of 0.5 mm and a kerf of 0.1 mm, i.e., a total length of 12.8 cm. The center frequency of the ultrasonic array is 3.5 MHz with over 80% bandwidth. PA signal is amplified 1500 times via self-built two-stage amplifier, and then received by a data acquisition system (Marsonics DAQ, Tianjin Langyuan Technology Co., Ltd. China) at 40 MHz sampling rate. The system is shown in Figure 6.
@@ 3.2.3 Parameters Selection and Computing Environment
In this vivo animal experiment, the specific values of the parameters we used are as follows:

### v2

@@ 3.2 In Vivo Human Experiment Results
Then in vivo human study data, including human arm was acquired by Shuang Li’s work[22] using synthetic matrix array. The system employed a nonfocusing linear array (customized by Imasonics, France) to receive PA signals. The linear array has 256 elements with a pitch of 0.5 mm and a kerf of 0.1 mm, i.e., a total length of 12.8 cm. The center frequency of the ultrasonic array is 3.5 MHz with over 80% bandwidth. PA signal is amplified 1500 times via self-built two-stage amplifier, and then received by a data acquisition system (Marsonics DAQ, Tianjin Langyuan Technology Co., Ltd. China) at 40 MHz sampling rate. The system is shown in Figure 7.
@@ 3.2.2 Comparison of Results before and after Optimization
In conclusion, in vivo human experiment, the reconstruction results obtained using the UBP algorithm and further optimized by the RI-SPPM algorithm also demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
@@ 3.2.3 Parameters Selection and Computing Environment
In this animal experiment, the specific values of the parameters we used are as follows: